OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (0, 0) and (3, 30). 10 0 5 −10

OpenStudy (anonymous):

if your two x values are 0 and 3, and your two y values are 0 and 30, set up the equation m=(y-y)/(x-x) Make sure the first y is above the first x, and the second y is above the second x. this gives you m=30-0/3-0 which =10

OpenStudy (nathan917):

When one of the points is (0,0), the slope formula becomes y/x and 30/3 = 10

OpenStudy (nathan917):

Hope i helped on the other answer.. its -14 i think.